OpenStudy (anonymous):

Find an equation of the line. Through (-5,-3); parallel to 2x+5y=7

OpenStudy (anonymous):

Set the equations 2x+5y=7 in y=mx+b form: 5y=-2x+7 y= (-2/5)x+7/5 The slope of the parallel line is -2/5 since parallel line have same slope. To find "b" substitute x and y in y=mx+b and solve -3=(-2/5)(-5)+b -3=2+b b=-5 Equation of parallel line passing through (-5,-3) and parallel to the line 2x+5y=7 is: y=(-2/5)x-5 or 5y=-2x-25 or 2x+5y=-25